# Poisoning Prevention in Federated Learning and Differential Privacy via Stateful Proofs of Execution

### v2

@@ V-B $\sf{SLAPP}$ in Detail
Protocol 1: $\sf{SLAPP}$ Protocol
Verifier ($\sf{\mathcal{V}rf}$)
  1. (1)
    Generate an authenticated $\sf{PoSX}$ request:
    $\sigma_{\sf{\mathcal{V}rf}}\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{V}rf}}}}},\textsf{H}({\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}}))$
  2. (2)
    Send ($\sigma_{\sf{\mathcal{V}rf}}$, $\sf{\mathcal{F}}$, $\sf{\mathcal{I}}$, $c_{\sf{\mathcal{V}rf}}$) to $\sf{\mathcal{P}rv}$
Prover ($\sf{\mathcal{P}rv}$)
In Non-Secure World:
  1. (3)
    Call Execute in Secure World with the received request
In Execute function, Secure World:
  1. (4)
    Authenticate the request and abort if $r=\bot$ or $exec=\top$:
    $r\leftarrow(c_{\sf{\mathcal{V}rf}}>c)\land\textsf{Verify}({\sf{pk_{{\sf{\mathcal{V}rf}}}}},\sigma_{\sf{\mathcal{V}rf}},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$
  2. (5)
    Update counter: $c\leftarrow c_{\sf{\mathcal{V}rf}}$ and initialize: $exec\leftarrow\top$, $stateUsed\leftarrow\bot$ $stateChecked\leftarrow\bot$
  3. (6)
    Disable interrupts and measure $PMEM$ and $\sf{\mathcal{V}rf}$ request:
    $h\leftarrow\textsf{H}(PMEM,{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}})$
  4. (7)
    Call ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ in Non-Secure World
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (8)
    Run ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$. Pass control to the Secure World when CheckState($s$) is called
In CheckState function, Secure World:
  1. (9)
    Perform an integrity check on $s$, store the result to $p$ and return to Non-Secure World: $stateChecked\leftarrow(\textsf{H}(s)\stackrel{{\scriptstyle?}}{{=}}{\sf{\mathcal{S}}}_{sec})$. Also, set $stateUsed\leftarrow\top$.
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (10)
    Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ execution. Pass control to the Secure World when SetState($s$) is called
In SetState function, Secure World:
  1. (11)
    Securely set ${\sf{\mathcal{S}}}_{sec}$ based on input $s$ and return to Non-Secure World:
    ${\sf{\mathcal{S}}}_{sec}\leftarrow\textsf{H}(s)\text{ if }(stateChecked\land exec)$
In $\sf{\mathcal{F}}$ function, Non-Secure World:
  1. (12)
    Continue with ${\sf{\mathcal{F}}}_{\sf{\mathcal{S}}}({\sf{\mathcal{I}}})$ until the execution is completed, producing output $\sf{\mathcal{O}}$, and then return to its caller, Execute, with $\sf{\mathcal{O}}$
In Execute function, Secure World:
  1. (13)
    Abort if $(exec~{}\land stateUsed~{}\land\neg stateChecked)$. Otherwise, include $\sf{\mathcal{O}}$ to the measurement and compute the proof:
    $\sigma\leftarrow\textsf{Sign}({\sf{sk_{{\sf{\mathcal{P}rv}}}}},\textsf{H}(h,{\sf{\mathcal{O}}}))$
  2. (14)
    Reset all flags: $exec\leftarrow\bot,stateChecked\leftarrow\bot,stateUsed\leftarrow\bot$
  3. (15)
    Enable interrupts and return with $({\sf{\mathcal{O}}},\sigma)$
In Non-Secure World:
  1. (16)
    Forward $({\sf{\mathcal{O}}},\sigma)$ to $\sf{\mathcal{V}rf}$
Verifier ($\sf{\mathcal{V}rf}$)
  1. (17)
    Increment $c_{\sf{\mathcal{V}rf}}$ and validate $\sigma$ by:
    $r\leftarrow\textsf{ValidatePoSX}({\sf{pk_{{\sf{\mathcal{P}rv}}}}},\sigma,PMEM^{\prime},{\sf{\mathcal{F}}},{\sf{\mathcal{I}}},c_{\sf{\mathcal{V}rf}},{\sf{\mathcal{O}}})$
    The protocol outputs $r$.
@@ VI $\sf{PoSX}$ Security Analysis
P3 - State Privacy. $\sf{SLAPP}$ protects privacy of $\sf{\mathcal{S}}$ from $\sf{\mathcal{V}rf}$ since the only information $\sf{\mathcal{V}rf}$ receives are $\sf{\mathcal{O}}$ and $\sigma$. In cases that $\sf{\mathcal{S}}$ can be leaked through $\sf{\mathcal{O}}$, $\sf{\mathcal{P}rv}$ has the option to refuse to send it over to $\sf{\mathcal{V}rf}$ by simply terminating the protocol. Moreover, since $\sigma$ is not a function of $\sf{\mathcal{S}}$, it leaks no information about $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$.
@@
Security Argument. We show that $\sf{SLAPP}$ satisfies both Definition IV and Definition IV implying adherence to Definition IV. Per Definition IV, IV-$\sf{PoX}$ security requires $\sf{PoX}$ assurance for stateless $\sf{\mathcal{F}}$ functions that run with input arguments, i.e., when ${\sf{\mathcal{S}}}=\phi$ and ${\sf{\mathcal{I}}}\neq\phi$. In this scenario, P2 and P5 ensure the atomic execution of $\sf{\mathcal{F}}$ with the authentic input $\sf{\mathcal{I}}$ provided by $\sf{\mathcal{V}rf}$, satisfying condition (i). P6 guarantees that $\sf{\mathcal{O}}$ is generated as a result of ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ execution, fulfilling condition (ii). With both conditions met, $\sf{SLAPP}$ achieves IV-$\sf{PoX}$ security. Meanwhile, SV-$\sf{PoX}$ security in Definition IV requires secure $\sf{PoX}$ of a stateful $\sf{\mathcal{F}}$ executed without input arguments. Similar to the previous argument, P5 and P6 directly address conditions (i) and (ii) of SV-$\sf{PoX}$ security even when ${\sf{\mathcal{I}}}=\phi$. As P3 ensures $\sf{\mathcal{S}}$ privacy, it fulfills condition (iii). Also, P4 guarantees that $\sf{\mathcal{S}}$ cannot be modified except by a fresh instance of $\sf{SLAPP}$ protocol, which satisfies condition (iv). Meeting all conditions in Definition IV, $\sf{SLAPP}$ is also SP-$\sf{PoX}$-secure. Lastly, $\sf{PoSX}$ security (per Definition IV) follows directly from simultaneous adherence to IV-$\sf{PoX}$ and SP-$\sf{PoX}$ security.∎
@@ VIII From $\sf{PoSX}$ to Poisoning Prevention
We now discuss how $\sf{SLAPP}$ can be leveraged to detect poisoning attacks. Unlike prior work, this guarantee extends to arbitrary functions $\sf{\mathcal{F}}$, whether they require private states or input arguments to execute. To illustrate this, we present ${}^{+}$ and ${}^{+}$, the poisoning-free versions of LDP-DC and FL-DC, respectively.
@@
${}^{+}$ utilizes $\sf{SLAPP}$ to ensure the correct execution of LDP phases:
@@
$\sf{SLAPP}$ in Setup phase ensures that the state variable $B$ is initialized to an empty value. During Collect phase, $\sf{SLAPP}$ protocol gives assurance to $\sf{\mathcal{V}rf}$ that: (1) $\sf{\mathcal{O}}$ is genuine, originating from a timely execution of the sensor function, and correctly privatized by the underlying LDP mechanism, (2) $B$ corresponded to the authentic value (e.g., empty in the first time of this phase’s execution) right before and during the protocol execution, and (3) $B$ is correctly updated as a result of the protocol execution. These prevent poisoning attacks because $\sf{\mathcal{A}dv}$ can tamper with neither $\sf{\mathcal{O}}$ (during Collect) nor $B$ (during Setup or Collect).
@@
${}^{+}$ follows a similar approach by using $\sf{SLAPP}$ to convey to $\sf{\mathcal{V}rf}$ the correct execution of:
@@
    Setup. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of Init-State function that initializes the local training dataset $D$ to an empty list on $\sf{\mathcal{P}rv}$. Abort if the protocol outputs $\bot$.
@@
Similar to ${}^{+}$, Setup guarantees to $\sf{\mathcal{V}rf}$ that $D$ starts empty. For Collect, $\sf{SLAPP}$ ensures that each record in $D$ is produced as a result of executing the expected sensor function on $\sf{\mathcal{P}rv}$. Finally, Train assures $\sf{\mathcal{V}rf}$ that the correct training function was applied to the authentic training data in $D$, leading to the received trained model $\sf{\mathcal{O}}$. In the context of FL, $\sf{\mathcal{A}dv}$ may attempt poisoning attacks in two ways: (1) model poisoning by directly tampering with the trained model $\sf{\mathcal{O}}$ and (2) data poisoning by manipulating the dataset $D$ used for training. Both are prevented by ${}^{+}$.
Remark. We emphasize that both ${}^{+}$ and ${}^{+}$ maintain the privacy of their original counterparts. The only additional information obtained by $\sf{\mathcal{V}rf}$ is $\sigma$, which is not a function of the underlying execution state. Also, as a system-level approach, these schemes offer a deterministic guarantee in differentiating poisoned data from authentic data, i.e., achieving 100% true positive and true negative rates, without relying on any assumptions about the underlying data.
@@ IX-A Experimental Setup
Cryptographic Variants: As noted in Section VII, $\sf{SLAPP}$ is agnostic to underlying cryptographic primitives. To showcase this flexibility, we provide 3 implementation variants of $\sf{SLAPP}$’s RoT: $\sf{SLAPP}_{SK}$, $\sf{SLAPP}_{PK}$, and $\sf{SLAPP}_{PQ}$:
@@
    $\sf{SLAPP}_{PQ}$ implements Sign and Verify using the quantum-resistant public-key signature Sphincs+, a low-RAM version of Sphincs-sha2-128f [^2] https://github.com/sphincs/low-ram-sphincsplus. This choice is supported by prior research [41, 26] showing feasibility of Sphincs+ on the target ARM Cortex-M devices.
Prototype. We prototype $\sf{SLAPP}$ variants on SAML11 Xplained Pro [36] development board as a representative of resource-constrained IoT devices. It features TrustZone-M on an ARM Cortex-M23 MCU @ 32MHz clock, with 64KB of FLASH (of which we assign 32KB to store the Non-secure World’s PMEM) and 16KB of RAM. To accurately isolate $\sf{SLAPP}$ base overhead, we implement a simple stateful $\sf{\mathcal{F}}$ function on $\sf{\mathcal{P}rv}$ that takes input from a GPIO port specified from a $\sf{PoSX}$ request, performs a sensor reading on that port, and outputs an accumulative sum of all readings over time to $\sf{\mathcal{V}rf}$. $\sf{\mathcal{V}rf}$ is deployed as a commodity desktop equipped with an Intel i7-8809G CPU @ 3.10GHz and 32GB RAM. $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ are connected via serial communication. All prototypes are open-sourced and publicly available at [3].
TABLE I: Lines of code and binary size of $\sf{SLAPP}$’s TCB. Overhead w.r.t. baseline is shown in parentheses.
Variants LoC Bin. Size in KB
$\sf{SLAPP}_{SK}$ 605 (+5.6%) 4.62 (+5.8%)
$\sf{SLAPP}_{PK}$ 1490 (+2.28%) 13.06 (+2.0%)
$\sf{SLAPP}_{PQ}$ 2806 (+1.24%) 13.28 (+2.1%)
@@
For comparison, we consider an alternative naive baseline approach in which all functions to prove execution $\sf{\mathcal{F}}$ are included as part of TCB residing within the Secure World. To perform $\sf{PoSX}$, TCB receives and authenticates a $\sf{PoSX}$ request from the Non-Secure World, just like $\sf{SLAPP}$. Unlike $\sf{SLAPP}$, this approach executes $\sf{\mathcal{F}}$ inside the Secure World, The result is signed (or MAC’ed) and forwarded to $\sf{\mathcal{V}rf}$, akin to $\sf{SLAPP}$. As this baseline approach is also agnostic to the underlying cryptographic primitives, we refer to $\sf{Baseline_{SK}}$, $\sf{Baseline_{PK}}$ and $\sf{Baseline_{PQ}}$ as the baseline approaches that utilize symmetric-key cryptography (HMAC-SHA256), public-key cryptography (ECDSA) and post-quantum cryptography (Sphincs+), respectively.
@@ IX-C Space Overhead
Code Size. $\sf{SLAPP}$’s TCB was implemented in C and compiled using the O3 optimization flag. Details of code size and the resulting binaries are shown in Table I. As $\sf{SLAPP}_{SK}$ relies on an inexpensive cryptographic primitive, it exhibits the smallest code size. Conversely, $\sf{SLAPP}_{PQ}$ results in the largest number of lines of C code. However, the O3 optimization was capable of reducing its binary size, making it comparable to that of $\sf{SLAPP}_{PK}$. Compared to the baseline, $\sf{SLAPP}$ introduces a small code size overhead, approximately 6% for $\sf{SLAPP}_{SK}$ and 2% for the others.
For $\sf{\mathcal{F}}$ instrumentation, $\sf{SLAPP}$ requires prepending a call to CheckState at the beginning of $\sf{\mathcal{F}}$ and another call to SetState before $\sf{\mathcal{F}}$ returns. This instrumentation results in only 2 additional lines of C code, enlarging $\sf{\mathcal{F}}$’s binary (residing in Non-Secure World) by a fixed 44 bytes.
Figure 4: Peak runtime data allocation across all approaches
Memory Usage. Next, we measure the peak amount of Secure World data allocated at runtime. This data consists of all stack and static/global variables (our implementation utilizes no heap allocation). As shown in Figure 4, $\sf{SLAPP}_{SK}$ and $\sf{SLAPP}_{PK}$ use roughly the same amount of data: $\approx$1.5KB. $\sf{SLAPP}_{PQ}$, on the other hand, requires more than double this amount: $\approx$3.6KB. Compared to the baseline, our $\sf{SLAPP}$ variants incur an additional memory usage ranging from 32 bytes for $\sf{SLAPP}_{PQ}$ to 200 bytes for $\sf{SLAPP}_{SK}$. These correspond to $<$ 1.25% of available RAM.
Figure 5: Execution time across all approaches

### v3

@@ VI $\sf{PoSX}$ Security Analysis
P3 - State Privacy. $\sf{SLAPP}$ protects privacy of $\sf{\mathcal{S}}$ from $\sf{\mathcal{V}rf}$ since the only information $\sf{\mathcal{V}rf}$ receives are $\sf{\mathcal{O}}$ and $\sigma$. $\sigma$ is not a function of $\sf{\mathcal{S}}$; thus it leaks nothing about $\sf{\mathcal{S}}$ to $\sf{\mathcal{V}rf}$. Thus, $\sf{SLAPP}$ incurs no leakage other than $\sf{\mathcal{O}}$ itself.
@@
Security Argument. We show that $\sf{SLAPP}$ satisfies both Definition IV and Definition IV implying adherence to Definition IV. Per Definition IV, IV-$\sf{PoX}$ security requires $\sf{PoX}$ assurance for stateless $\sf{\mathcal{F}}$ functions that run with input arguments, i.e., when ${\sf{\mathcal{S}}}=\phi$ and ${\sf{\mathcal{I}}}\neq\phi$. In this scenario, P2 and P5 ensure the atomic execution of $\sf{\mathcal{F}}$ with the authentic input $\sf{\mathcal{I}}$ provided by $\sf{\mathcal{V}rf}$, satisfying condition (i). P6 guarantees that $\sf{\mathcal{O}}$ is generated as a result of ${\sf{\mathcal{F}}}({\sf{\mathcal{I}}})$ execution, fulfilling condition (ii). With both conditions met, $\sf{SLAPP}$ achieves IV-$\sf{PoX}$ security. Meanwhile, SV-$\sf{PoX}$ security in Definition IV requires secure $\sf{PoX}$ of a stateful $\sf{\mathcal{F}}$ executed without input arguments. Similar to the previous argument, P5 and P6 directly address conditions (i) and (ii) of SV-$\sf{PoX}$ security even when ${\sf{\mathcal{I}}}=\phi$. As P3 ensures that $\sf{SLAPP}$ leaks nothing about $\sf{\mathcal{S}}$ besides its intended execution output, it fulfills condition (iii). Also, P4 guarantees that $\sf{\mathcal{S}}$ cannot be modified except by a fresh instance of $\sf{SLAPP}$ protocol, which satisfies condition (iv). Meeting all conditions in Definition IV, $\sf{SLAPP}$ is also SP-$\sf{PoX}$-secure. Lastly, $\sf{PoSX}$ security (per Definition IV) follows directly from simultaneous adherence to IV-$\sf{PoX}$ and SP-$\sf{PoX}$ security.∎
Remark. $\sf{SLAPP}$ maintains P1-P6 even in the presence of an adaptive $\sf{\mathcal{A}dv}$; this implies that irrespective of $\sf{\mathcal{A}dv}$ actions or knowledge of $\sf{SLAPP}$ (assuming no invasive hardware attacks), none of these properties can be compromised.
@@ VIII From $\sf{PoSX}$ to Poisoning Prevention
We now discuss how $\sf{SLAPP}$ can be leveraged to detect poisoning attacks in LDP-DC and FL-DC.
@@
LDP-DC+ utilizes $\sf{SLAPP}$ to ensure the correct execution of LDP steps:
@@
$\sf{SLAPP}$ in Setup step ensures that the state variable $B$ is initialized to an empty value. At a later time, Collect can be executed, where $\sf{SLAPP}$ protocol gives assurance to $\sf{\mathcal{V}rf}$ that: (1) $\sf{\mathcal{O}}$ is genuine, originating from a timely execution of the sensor function, and correctly privatized by the underlying LDP mechanism, (2) $B$ corresponded to the authentic value (e.g., empty at the first time of this step’s execution) right before and during the protocol execution, and (3) $B$ is correctly updated as a result of the protocol execution. These prevent poisoning attacks because $\sf{\mathcal{A}dv}$ can tamper with neither $\sf{\mathcal{O}}$ (during Collect) nor $B$ (during Setup or Collect).
@@
FL-DC+ follows a similar approach by using $\sf{SLAPP}$ to convey to $\sf{\mathcal{V}rf}$ the correct execution of:
@@
    Setup. Run $\sf{SLAPP}$ protocol to obtain a $\sf{PoSX}$ of Init-State function that sets the local training dataset $D$ to an empty list on $\sf{\mathcal{P}rv}$. Abort if it outputs $\bot$.
@@
Similar to LDP-DC+, Setup guarantees to $\sf{\mathcal{V}rf}$ that $D$ starts empty. For Collect, $\sf{SLAPP}$ ensures that each record in $D$ is produced as a result of executing the expected sensor function on $\sf{\mathcal{P}rv}$. Finally, Train assures $\sf{\mathcal{V}rf}$ that the correct training function was applied to the authentic training data in $D$, leading to the received trained model $\sf{\mathcal{O}}$. In the context of FL, $\sf{\mathcal{A}dv}$ may attempt poisoning attacks in two ways: (1) model poisoning by directly tampering with the trained model $\sf{\mathcal{O}}$ and (2) data poisoning by manipulating the dataset $D$ used for training. Both are prevented by LDP-DC+.
Remark. We emphasize that both LDP-DC+ and FL-DC+ maintain the privacy of their original counterparts. The only additional information besides $\sf{\mathcal{O}}$ obtained by $\sf{\mathcal{V}rf}$ is $\sigma$, which is not a function of the underlying execution state. Also, as a system-level approach, these schemes offer a deterministic guarantee in discerning poisoned data from authentic data, i.e., achieving 100% true positive and true negative rates, irrespective of adaptive attacks. It also comes without any assumptions about the data distributions on $\sf{\mathcal{P}rv}$.
@@ IX-A Experimental Setup
Cryptographic Variants. As noted in Section VII, $\sf{SLAPP}$ is agnostic to underlying cryptographic primitives. To showcase this flexibility, we provide 3 implementation variants of $\sf{SLAPP}$’s RoT: $\sf{SLAPP}_{SK}$, $\sf{SLAPP}_{PK}$, and $\sf{SLAPP}_{PQ}$:
@@
    $\sf{SLAPP}_{PQ}$ implements Sign and Verify using the quantum-resistant public-key signature Sphincs+, a low-RAM version of Sphincs-sha2-128f [^2] https://github.com/sphincs/low-ram-sphincsplus. This choice is supported by prior research [41, 27] showing feasibility of Sphincs+ on Cortex-M devices.
Prototype. We prototype $\sf{SLAPP}$ variants on a NUCLEO-L552ZE-Q [51] development board, representing resource-constrained IoT devices. It features TrustZone-M on an ARM Cortex-M33 MCU @ 110MHz clock, with 512KB of FLASH (of which we assign 256KB to store the Non-secure World’s PMEM) and 256KB of RAM. To accurately isolate $\sf{SLAPP}$ overheads, we implement a simple stateful $\sf{\mathcal{F}}$ function on $\sf{\mathcal{P}rv}$ that takes input from a GPIO port specified from a $\sf{PoSX}$ request, performs a sensor reading on that port, and outputs an accumulative sum of all readings over time to $\sf{\mathcal{V}rf}$. $\sf{\mathcal{V}rf}$ is deployed as a commodity desktop equipped with an Intel i5-9300H CPU @ 2.4GHz. $\sf{\mathcal{V}rf}$ and $\sf{\mathcal{P}rv}$ are connected via serial communication. All prototypes are open-sourced and publicly available at [47].
@@
For comparison, we consider an alternative naive baseline approach in which all functions to prove execution $\sf{\mathcal{F}}$ are included as part of TCB in the Secure World. To perform $\sf{PoSX}$, TCB receives and authenticates a request from the Non-Secure World, just like $\sf{SLAPP}$. Unlike $\sf{SLAPP}$, this approach executes $\sf{\mathcal{F}}$ inside the Secure World, The result is signed (or MAC’ed) and forwarded to $\sf{\mathcal{V}rf}$, akin to $\sf{SLAPP}$. As this baseline approach is also agnostic to the underlying cryptographic primitives, we refer to $\sf{Baseline_{SK}}$, $\sf{Baseline_{PK}}$ and $\sf{Baseline_{PQ}}$ as the baseline approaches that utilize symmetric-key cryptography (HMAC-SHA256), public-key cryptography (ECDSA) and post-quantum cryptography (Sphincs+), respectively.
@@
TABLE I: Binary size (in KB) of TCB.
Variants $\sf{Baseline_{SK}}$ $\sf{SLAPP}$ Overhead
$\sf{SLAPP}_{SK}$ 17.0 17.5 2.9%
$\sf{SLAPP}_{PK}$ 34.5 35.0 1.4%
$\sf{SLAPP}_{PQ}$ 24.0 24.5 2.0%
@@ IX-C Space Overhead
Code Size. $\sf{SLAPP}$’s TCB was implemented in C and compiled using the -O3 optimization flag. Details of code size are shown in Table I. As $\sf{SLAPP}_{SK}$ relies on an inexpensive cryptographic primitive, it exhibits the smallest code size. Conversely, $\sf{SLAPP}_{PK}$ results in the largest binary size of 35.0KB while $\sf{SLAPP}_{PQ}$’s binary is around 24.5KB. Compared to the baselines, $\sf{SLAPP}$ introduces a small code size overhead (0.5KB), corresponding to 1.4-2.9% across all variants.
For $\sf{\mathcal{F}}$ instrumentation, $\sf{SLAPP}$ requires prepending a call to CheckState at the beginning of $\sf{\mathcal{F}}$ and another call to SetState before $\sf{\mathcal{F}}$ returns. This instrumentation results in only 2 additional lines of C code, enlarging $\sf{\mathcal{F}}$’s binary (residing in Non-Secure World) by a fixed 22 bytes.
(a) Peak runtime data allocation
(b) Average execution time
Figure 3: Resource usage across all approaches
Memory Usage. Next, we estimate the peak amount of Secure World data allocated at runtime. This data consists of all stack and static/global variables (our implementation utilizes no heap allocation). As shown in Fig. 3(a), $\sf{SLAPP}_{SK}$ and $\sf{SLAPP}_{PK}$ uses roughly the same amount of data: 2.74 and 3.29KB. $\sf{SLAPP}_{PQ}$, on the other hand, requires more than triple this amount: 9.22KB. Compared to the baselines, our $\sf{SLAPP}$ variants incur an additional memory usage ranging from 50 bytes for $\sf{SLAPP}_{SK}$ to 2.38KB for $\sf{SLAPP}_{PQ}$. These correspond to $<$ 4% of available RAM in the Secure World.